OpenStudy (anonymous):

The first four terms of a sequence are shown below. 8, 5, 2, -1 Which of the following functions best defines this sequence?

OpenStudy (anonymous):

please help!

OpenStudy (solomonzelman):

you said ` which of the following `, so can you please post the options ?

OpenStudy (anonymous):

1) f(1) = 8, f(n + 1) = f(n) + 3; for n ≥ 1 2) f(1) = 8, f(n + 1) = f(n) - 5; for n ≥ 1 3) f(1) = 8, f(n + 1) = f(n) + 5; for n ≥ 1 4) f(1) = 8, f(n + 1) = f(n) - 3; for n ≥ 1

OpenStudy (solomonzelman):

The recursive formula for this sequence is \(\Large\color{black}{ f_{n+1}=f_{n}-3 }\). In other words, to get to the next term, you subtract 3 from the term before it. And the first term is 8. So the answer is ?

OpenStudy (solomonzelman):

If you need more help, tell me....

OpenStudy (anonymous):

I don't understand